OpenStudy (anonymous):

What is the equation of the line that passes through the point (4, 1) and is perpendicular to the line y = –2x + 4?

hero (hero):

Well, if it's perpendicular to the given line, it will have a slope of 1/2, so, we can use the point-slope formula to find it: y - y_1 = m(x -x_1) y - 1 = 1/2(x - 4) y - 1 = x/2 - 2 y = x/2 -1

OpenStudy (anonymous):

y = x/2 + b 1 = 4/2 + b b = -1 y = x/2 -1

hero (hero):

Money bird used y = mx +b to solve for b. That's the alternative way to do it.